# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v1

@@ 5.3 Realistic Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
Method PDM-Lite [69] PlanT [71]
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ BR $\downarrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ BR $\downarrow$
Pluto [53] 77.84 $\pm$ 2.20 81.84 $\pm$ 2.20 0.96 $\pm$ 0.00 23.33 $\pm$ 5.77 42.52 $\pm$ 4.72 52.53 $\pm$ 5.29 0.79 $\pm$ 0.01 46.67 $\pm$ 15.28
PPO[14] 76.26 $\pm$ 0.12 76.26 $\pm$ 0.12 1.00 $\pm$ 0.00 30.00 $\pm$ 0.00 36.39 $\pm$ 1.11 58.30 $\pm$ 3.41 0.66 $\pm$ 0.05 43.33 $\pm$ 11.55
FREA[14] 83.53 $\pm$ 0.13 83.53 $\pm$ 0.13 1.00 $\pm$ 0.00 20.00 $\pm$ 0.00 39.61 $\pm$ 1.34 53.82 $\pm$ 2.20 0.73 $\pm$ 0.01 56.67 $\pm$ 5.77
FPPO-RS[14] 83.52 $\pm$ 0.09 83.52 $\pm$ 0.09 1.00 $\pm$ 0.00 20.00 $\pm$ 0.00 38.85 $\pm$ 4.91 52.27 $\pm$ 5.38 0.66 $\pm$ 0.07 50.00 $\pm$ 10.00
SFT-Pluto 86.09 $\pm$ 2.04 90.16 $\pm$ 1.82 0.97 $\pm$ 0.03 13.33 $\pm$ 5.77 39.41 $\pm$ 4.97 57.90 $\pm$ 3.72 0.70 $\pm$ 0.06 50.00 $\pm$ 0.00
RS-Pluto [22] 89.32 $\pm$ 1.41 89.32 $\pm$ 1.41 1.00 $\pm$ 0.00 13.33 $\pm$ 5.77 42.05 $\pm$ 4.08 59.58 $\pm$ 3.27 0.66 $\pm$ 0.02 50.00 $\pm$ 0.00
RTR-Pluto [21] 87.64 $\pm$ 1.56 91.64 $\pm$ 1.56 0.96 $\pm$ 0.00 10.00 $\pm$ 0.00 40.08 $\pm$ 2.38 55.87 $\pm$ 1.71 0.69 $\pm$ 0.04 50.00 $\pm$ 0.00
PPO-Pluto 85.63 $\pm$ 2.02 89.63 $\pm$ 2.02 0.96 $\pm$ 0.00 16.67 $\pm$ 5.77 41.86 $\pm$ 2.78 59.82 $\pm$ 2.90 0.66 $\pm$ 0.04 43.33 $\pm$ 5.77
REINFORCE-Pluto 92.17 $\pm$ 3.45 92.17 $\pm$ 3.45 1.00 $\pm$ 0.00 10.00 $\pm$ 10.00 45.25 $\pm$ 1.75 63.74 $\pm$ 5.91 0.67 $\pm$ 0.05 43.33 $\pm$ 5.77
GRPO-Pluto 89.86 $\pm$ 2.10 91.19 $\pm$ 3.46 0.99 $\pm$ 0.02 6.67 $\pm$ 5.77 47.24 $\pm$ 5.67 61.88 $\pm$ 0.71 0.75 $\pm$ 0.07 $3$$\pm$ 5.77
RIFT-Pluto (ours) 94.78 $\pm$ 1.37 94.78 $\pm$ 1.37 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 44.28 $\pm$ 3.15 61.94 $\pm$ 5.08 0.64 $\pm$ 0.02 40.00 $\pm$ 10.00
Table 3: Ablation study on RIFT. Evaluation under PDM-Lite [69] AV setting with three random seeds.
Method Controllability Realism
Infraction Progress Safety Speed Acceleration
ORR $\downarrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ SW $\uparrow$ WD $\downarrow$ SW $\uparrow$
w/ Aggressive 0.91 (+0.55) 8.41 (+1.58) 1053.76 (+58.43) 2.93 (+0.19) 2.88 (+0.17) 0.97 (+0.00) 3.89 (-0.57) 0.94 (+0.01)
w/ All-head 0.43 (+0.07) 7.84 (+1.01) 827.12 (168.21) 2.83 (+0.09) 2.76 (+0.05) 0.96 (-0.01) 4.70 (+0.24) 0.94 (+0.01)
w/ PPO-Clip 0.04 (-0.32) 2.03 (-4.80) 655.39 (-339.94) 2.57 (-0.17) 2.54 (-0.17) 0.91 (-0.06) 5.92 (+1.46) 0.94 (+0.01)
RIFT-Pluto (ours) 0.36 6.83 995.33 2.74 2.71 0.97 4.46 0.93
@@ 5.4 Ablation Study (Q3)
In exploring the design choices within RIFT, we conduct a series of ablation studies focusing on three key aspects: the choice of policy optimization clip (Dual-Clip vs. PPO-Clip), the fine-tuning module (Trajectory Scoring Head vs. All Head), and the reward design (Normal vs. Aggressive). All experiments are conducted under consistent settings, and results are summarized in Table 3.
Dual-Clip vs. PPO-Clip. We first evaluate the effect of the dual-clip mechanism by replacing it with the standard PPO-Clip (w/ PPO-Clip). This variant leads to more conservative and less realistic behaviors, as evidenced by reduced Progress and degraded Speed Realism metrics.
Trajectory Scoring Head vs. All Head. Next, we examine the impact of the fine-tuning module on performance by comparing RIFT fine-tuned using only the trajectory scoring head with a variant fine-tuned on all heads (w/ All Head). The latter results in a slight drop across Realism, Progress, and Infraction metrics, likely due to the disruption of trajectory-level realism captured during pre-training.
Normal vs. Aggressive. Finally, we explore the effect of different reward preferences on controllability by adopting a more aggressive reward focused on driving efficiency (w/ Aggressive). This variant improves Progress and Speed Realism but increases Infractions, reflecting the intended trade-off and confirming that RIFT allows flexible reward design while preserving user controllability. Additional qualitative results regarding controllability are presented in Appendix D.1.
@@ 6 Conclusion
In this work, we propose a dual-stage AV-centered simulation framework that conducts open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and multimodality, followed by closed-loop RL fine-tuning in a physics-based simulator to address the covariate shift problem and enhance controllability. During fine-tuning, we introduce RIFT, a simple yet effective closed-loop RL fine-tuning strategy that preserves the trajectory-level multimodality through a GRPO-style group-relative advantage formulation and enhances both controllability and training stability by replacing KL regularization with the dual-clip mechanism. Extensive experiments demonstrate that RIFT generates high-quality traffic scenarios that are both realistic and controllable, offering a reliable simulation environment for evaluating AV performance in diverse and interactive scenarios, thus bridging the gap between scenario generation and closed-loop AV evaluation.
Limitations and Future Work. Currently, RIFT adopts a simple state-wise reward model, $\mathrm{StateWiseRM}$, as detailed in Section B.6, to evaluate each trajectory point without forward simulation. This may result in inaccurate trajectory-level reward estimations over time, potentially affecting closed-loop RL fine-tuning performance (further discussed in Section D.2). Additionally, the current implementation of our traffic simulation framework supports only AV methods with fully observable, vectorized scene representations. Extending this framework to support end-to-end closed-loop training and evaluation remains a promising direction for future research.
@@ A.1 Problem Setup
For a given state $s$, let $\pi_{\theta}(\tau\mid s)$ denote the probability of trajectory $\tau$ under policy parameters $\theta$. Define the success probability
@@
The (group-)relative advantage [20] in Equation 6 therefore reduces to
@@ A.5 Detailed Proofs
Throughout this subsection, we reuse the notation introduced in Section A.4:
@@
Proof of Proposition A.3.
Define $g(p)=h_{\epsilon}(p)-p$. Lemma A.2 gives $g(0)>0$ and $g(1)<0$, while Lemma A.2 also guarantees the continuity of $g$. The Intermediate Value Theorem therefore yields a root $p^{\star}\in(0,1)$ with $p^{\star}=h_{\epsilon}(p^{\star})$. ∎
Proof of Theorem A.4.
@@
Fix $\delta\in(0,\tfrac{1}{2})$ such that the first iterate satisfies $h_{\epsilon}(p_{0})\in[\delta,1-\delta]$ (Lemma A.2 ensures such a $\delta$ exists). Then $|\Phi^{\prime}(p)|\leq L_{\delta}(\epsilon):=\epsilon/\sqrt{\delta(1-\delta)}$ for every subsequent $p$.
@@ B.1 Experiment Framework
To enable reliable AV-centered closed-loop simulation, we develop our framework based on widely adopted closed-loop traffic simulation platforms, such as the CARLA Leaderboard [70] and Bench2Drive [46]. Traditionally, these platforms use predefined scenarios along the AV’s global route to evaluate the multi-dimensional performance of AV methods. In contrast, we replace these static scenarios with dynamically generated traffic flows by randomly spawning background vehicles around the AV’s global path and simulating their behavior using rule-based driving policies, as described in Section 3.1. Through the CBV identification mechanism outlined in Section B.2, we naturally introduce interactions between the AV and CBVs, thereby generating continuous, interactive scenarios over time. This framework serves as the foundation for both the training and evaluation processes in this paper.
B.2 CBV Identification Details
Identifying Critical Background Vehicles (CBVs) is essential to our AV-centered closed-loop simulation. Let $\mathcal{V}_{\text{AV}}$ denote the autonomous vehicle (AV), and $\mathcal{V}_{\text{BV}}=\{\mathcal{V}_{i}\}_{i=1}^{N}$ represent the set of background vehicles in the environment. The AV navigates along a predefined global route $\mathcal{P}=\{p_{k}\}_{k=1}^{M}$, where each $p_{k}$ corresponds to a waypoint along the route. The goal of CBV identification is to select background vehicles that are likely to share the AV’s destination and have similar estimated travel distance, thereby facilitating route-level interactions between the AV and CBVs. The primary criterion for identifying CBVs is the relative distance-to-goal difference between the AV and each background vehicle. This is mathematically expressed as:
@@ B.2 CBV Identification Details
where, $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})$ and $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})$ denote the estimated travel distance required for the background vehicle $\mathcal{V}_{i}$ and the AV to reach waypoint $p_{k}$, respectively. The distance-to-goal for each vehicle is computed by determining the distance from its current position to the target waypoint $p_{k}$ using the A* global path planning algorithm [72]. A threshold $\delta$ is introduced to define the maximum allowable difference in distance-to-goal. A background vehicle is considered critical and included in the CBV set $\mathcal{C}$ if the absolute distance-to-goal difference between it and the AV is smaller than $\delta$.
This approach selects background vehicles whose destinations and estimated travel distances are sufficiently aligned with those of the AV, thereby ensuring meaningful and realistic route-level interactions. Once a CBV is identified, the planning path previously generated via A* during distance-to-goal estimation is directly adopted as its global navigation path, naturally introducing route-level interactions between the AV and CBVs. The threshold $\delta$ serves as a tunable parameter to adjust the sensitivity of the CBV selection process. In this study, we set $\delta$ to $15m$ to achieve a balanced trade-off between sensitivity and selection accuracy.
@@ B.3 Algorithm Framework
For clarity, we summarize the procedure of RIFT within our AV-centered closed-loop simulation framework in Algorithm 1. The trajectory generation model initializes from the IL pre-trained checkpoint provided by Pluto official codebase [^1] https://github.com/jchengai/pluto, followed by closed-loop RL fine-tuning within the CARLA simulator [62] to generate realistic and controllable traffic scenarios.
Algorithm 1 Procedure for RIFT in the AV-Centered Closed-Loop Simulation Framework.
1:Input: IL pre-trained trajectory generation model $\pi_{\theta_{\text{init}}}$, buffer $\mathcal{D}$triangleright IL pre-training (nuPlan [11])
2:trajectory generation model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$
3:for $\text{iteraction}=1,\dots,I$ do $\triangleright$ Closed-loop RL fine-tuning (CARLA [62])
4:     Update the old trajectory generation model $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
@@
Compute reward $\{r_{i}\}_{i=1}^{G}$, advantage $\{\hat{A}_{i}\}_{i=1}^{G}$ for each trajectory $\tau_{i}$ with Equation 6
@@
10:         end for
11:     end while
12:     for RIFT $\text{iteration}=1,\dots,\mu$ do $\triangleright$ Policy fine-tuning
13:
@@
14:
Update model $\pi_{\theta}$ by maximizing the RIFT objective (Equation 8)
15:     end for
@@
17:Output: RL fine-tuned trajectory generation model

### v2

@@ 5.3 Generated Traffic Scenarios for Closed-Loop AV Evaluation (Q2)
Method PDM-Lite Beibwenger2024pdmLite PlanT Renz2022PlanT UniAD hu2023_uniad VAD jiang2023vad
DS $\uparrow$ BR $\uparrow$ DS $\Delta{\text{DS}}\downarrow$ DS $\Delta{\text{DS}}\downarrow$ DS $\Delta{\text{DS}}\downarrow$
Pluto cheng2024pluto 77.84 $\pm$ 2.20 23.33 $\pm$ 5.77 42.52 $\pm$ 4.72 -35.32 73.73 $\pm$ 1.24 -4.11 66.87 $\pm$ 2.11 -10.97
PPOchen2024frea 76.26 $\pm$ 0.12 30.00 $\pm$ 0.00 36.39 $\pm$ 1.11 -39.87 69.79 $\pm$ 1.41 -6.47 67.64 $\pm$ 1.27 -8.62
FREAchen2024frea 83.53 $\pm$ 0.13 20.00 $\pm$ 0.00 39.61 $\pm$ 1.34 -43.92 69.29 $\pm$ 5.22 -14.24 67.57 $\pm$ 5.37 -15.96
FPPO-RSchen2024frea 83.52 $\pm$ 0.09 20.00 $\pm$ 0.00 38.85 $\pm$ 4.91 -44.67 75.13 $\pm$ 5.18 -8.39 69.15 $\pm$ 2.79 -14.37
SFT-Pluto 86.09 $\pm$ 2.04 13.33 $\pm$ 5.77 39.41 $\pm$ 4.97 -47.28 77.49 $\pm$ 5.93 -9.20 68.89 $\pm$ 0.87 -17.80
RS-Pluto peng2024improving 89.32 $\pm$ 1.41 13.33 $\pm$ 5.77 42.05 $\pm$ 4.08 -47.27 80.62 $\pm$ 0.78 -8.70 69.48 $\pm$ 5.02 -19.84
RTR-Pluto zhang2023rtr 87.64 $\pm$ 1.56 10.00 $\pm$ 0.00 40.08 $\pm$ 2.38 -47.56 77.69 $\pm$ 2.82 -9.95 66.27 $\pm$ 4.53 -21.37
PPO-Pluto 85.63 $\pm$ 2.02 16.67 $\pm$ 5.77 41.86 $\pm$ 2.78 -43.77 77.14 $\pm$ 3.36 -8.49 68.62 $\pm$ 3.16 -17.01
REINFORCE-Pluto 92.17 $\pm$ 3.45 10.00 $\pm$ 10.00 45.25 $\pm$ 1.75 -46.92 79.89 $\pm$ 1.97 -12.28 70.28 $\pm$ 3.58 -21.89
GRPO-Pluto 89.86 $\pm$ 2.10 6.67 $\pm$ 5.77 47.24 $\pm$ 5.67 -42.62 81.02 $\pm$ 0.64 -8.84 72.55 $\pm$ 0.74 -17.31
RIFT-Pluto (ours) 94.78 $\pm$ 1.37 0.00 $\pm$ 0.00 44.28 $\pm$ 3.15 -50.50 73.79 $\pm$ 6.53 -20.99 68.24 $\pm$ 3.23 -26.54
Table 3: Ablation Study on RIFT. Evaluation under PDM-Lite Beibwenger2024pdmLite AV setting with three random seeds.
Method Kinematic Metrics Interaction Metrics Map Metrics
S-SW $\uparrow$ S-WD $\downarrow$ A-SW $\uparrow$ CPK $\downarrow$ RP $\uparrow$ 2D-TTC $\uparrow$ ACT $\uparrow$ ORR $\downarrow$
w/ Aggressive 0.97 (+0.00) 3.89 (-0.57) 0.94 (+0.01) 8.41 (+1.58) 1053.76 (+58.43) 2.93 (+0.19) 2.88 (+0.17) 0.91 (+0.55)
w/ KL 0.93 (-0.04) 5.33 (+0.87) 0.90 (-0.03) 7.05 (+0.22) 815.06 (-180.27) 2.76 (+0.02) 2.73 (+0.02) 0.38 (+0.02)
w/ All-head 0.96 (-0.01) 4.70 (+0.24) 0.94 (+0.01) 7.84 (+1.01) 827.12 (-168.21) 2.83 (+0.09) 2.76 (+0.05) 0.43 (+0.07)
w/ PPO-Clip 0.91 (-0.06) 5.92 (+1.46) 0.94 (+0.01) 2.03 (-4.80) 655.39 (-339.94) 2.57 (-0.17) 2.54 (-0.17) 0.04 (-0.32)
RIFT-Pluto (ours) 0.97 4.46 0.93 6.83 995.33 2.74 2.71 0.36
@@ 5.4 Ablation Study (Q3)
In exploring the design choices within RIFT, we conduct a series of ablation studies focusing on four key aspects: the choice of policy optimization clip (Dual-Clip vs. PPO-Clip), the fine-tuning module (Scoring Head vs. All Head), the choice of KL regularization (w/ KL vs. w/o KL), and the reward design (Normal vs. Aggressive). All experiments are conducted under consistent settings, and results are summarized in Table˜3.
Dual-Clip vs. PPO-Clip. We first evaluate the effect of the dual-clip mechanism by replacing it with the standard PPO-Clip (w/ PPO-Clip). This variant leads to more conservative and less realistic behaviors, as evidenced by reduced driving efficiency and degraded kinematic metrics.
Scoring Head vs. All Head. Next, we examine the impact of the fine-tuning module on performance by comparing RIFT fine-tuned using only the scoring head with a variant fine-tuned on all heads (w/ All Head). The latter induces a slight degradation across metrics, primarily due to the disruption of trajectory-level realism learned during pre-training.
w/ KL vs. w/o KL. Then, we ablate the KL term that constrains divergence from the IL pre-trained policy. As shown in Table˜3, removing this constraint improves controllability without degrading realism. This validates our decoupled design: instead of enforcing distributional consistency via KL, we freeze the generation head and fine-tune only the scoring head. This separation enables flexible trajectory selection while preserving the trajectory-level realism inherited from pre-training.
Normal vs. Aggressive. Finally, we explore the effect of different reward preferences on controllability by adopting a more aggressive reward focused on driving efficiency (w/ Aggressive). This variant improves driving efficiency but increases collision frequency and off-road rate, highlighting the intended trade-off and demonstrating that RIFT supports flexible reward design while maintaining style-level controllability. Additional qualitative results regarding controllability are presented in Appendix D.1.
@@ 6 Conclusion
In this work, we propose a dual-stage AV-centric simulation framework that conducts open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and route-level controllability, followed by closed-loop RL fine-tuning in a physics-based simulator to address covariate shift and enhance style-level controllability. During fine-tuning, we introduce RIFT, a novel RL fine-tuning strategy that evaluates all candidate modalities using group-relative optimization combined with a dual-clip surrogate objective, thereby enhancing style-level controllability and mitigating covariate shift, while preserving the trajectory-level realism and route-level controllability established in IL pre-training. Extensive experiments demonstrate that RIFT generates scenarios with superior realism and controllability, effectively revealing the limitations of modern AV systems and further bridging the gap between traffic simulation and reliable closed-loop evaluation.
Limitations and Future Work. In the current framework, the generation head is frozen during closed-loop fine-tuning, and its reliability solely stems from the robust trajectory generation capability learned during IL pre-training. However, without reliable expert demonstrations, the realism and robustness of generated trajectories cannot be further improved during closed-loop fine-tuning. This limitation highlights a key avenue for future research: developing methods—potentially leveraging RL or other self-improvement paradigms—that can enhance the quality of trajectory generation without relying on expert demonstrations.
@@ A.1 Problem Setup
For a given state $s$, let $\pi_{\theta}(\tau\mid s)$ denote the probability of trajectory $\tau$ under policy parameters $\theta$. For simplicity, we use $\tau$ to denote both the rollout and the candidate trajectory, as there exists a fixed mapping between them. Define the success probability
@@
The (group-)relative advantage shao2024grpo in Equation˜6 therefore reduces to
@@ A.5 Detailed Proofs
Throughout this subsection, we reuse the notation introduced in Section˜A.4:
@@
Proof of Proposition˜A.3.
Define $g(p)=h_{\epsilon}(p)-p$. Lemma˜A.2 gives $g(0)>0$ and $g(1)<0$, while Lemma˜A.2 also guarantees the continuity of $g$. The Intermediate Value Theorem therefore yields a root $p^{\star}\in(0,1)$ with $p^{\star}=h_{\epsilon}(p^{\star})$. ∎
Proof of Theorem˜A.4.
@@
Fix $\delta\in(0,\tfrac{1}{2})$ such that the first iterate satisfies $h_{\epsilon}(p_{0})\in[\delta,1-\delta]$ (Lemma˜A.2 ensures such a $\delta$ exists). Then $|\Phi^{\prime}(p)|\leq L_{\delta}(\epsilon):=\epsilon/\sqrt{\delta(1-\delta)}$ for every subsequent $p$.
@@ B.1 Experiment Framework
Our framework for reliable AV-centric closed-loop simulation is developed upon well-established traffic simulation platforms, notably the CARLA Leaderboard CARLA_Leaderboard and Bench2Drive jia2024bench2drive , which serve as standard benchmarks in autonomous driving research. Traditionally, these platforms use predefined scenarios along the AV’s global route to evaluate the multi-dimensional performance of AV methods. In contrast, we replace these static scenarios with dynamically generated traffic flows by randomly spawning background vehicles around the AV’s global path and simulating their behavior using rule-based driving policies, as described in Section˜3.1. Through the CBV identification mechanism outlined in Section˜B.2, we naturally introduce interactions between the AV and CBVs, thereby generating continuous, interactive scenarios over time. This framework serves as the foundation for both the training and evaluation processes in this paper.
B.2 Route-level Analysis for CBV Identification
Identifying Critical Background Vehicles (CBVs) is essential to our AV-centric closed-loop simulation. Let $\mathcal{V}_{\text{AV}}$ denote the autonomous vehicle (AV), and $\mathcal{V}_{\text{BV}}=\{\mathcal{V}_{i}\}_{i=1}^{N}$ represent the set of background vehicles in the environment. The AV navigates along a predefined global route $\mathcal{P}=\{p_{k}\}_{k=1}^{M}$, where each $p_{k}$ corresponds to a waypoint along the route. The goal of CBV identification is to select background vehicles that are likely to share the AV’s destination and have similar estimated travel distance, thereby facilitating route-level interactions between the AV and CBVs. The primary criterion for identifying CBVs is the relative distance-to-goal difference between the AV and each background vehicle. This is mathematically expressed as:
@@ B.2 Route-level Analysis for CBV Identification
where, $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})$ and $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})$ denote the estimated travel distance required for the background vehicle $\mathcal{V}_{i}$ and the AV to reach waypoint $p_{k}$, respectively. The distance-to-goal for each vehicle is computed by determining the distance from its current position to the target waypoint $p_{k}$ using the A* global path planning algorithm hart1968a_star . A threshold $\delta$ is introduced to define the maximum allowable difference in distance-to-goal. A background vehicle is considered critical and included in the CBV set $\mathcal{C}$ if the absolute distance-to-goal difference between it and the AV is smaller than $\delta$.
This approach selects background vehicles whose destinations and estimated travel distances are sufficiently aligned with those of the AV, thereby ensuring meaningful and realistic route-level interactions. Once a CBV is identified, the planning path previously generated via A* during distance-to-goal estimation is directly adopted as its global navigation path, which is further transformed into the reference line for downstream CBV planning, naturally introducing route-level interactions between the AV and CBVs. The threshold $\delta$ serves as a tunable parameter to adjust the sensitivity of the CBV selection process. In this study, we set $\delta$ to $15m$ to achieve a balanced trade-off between sensitivity and selection accuracy.
@@ B.3 Algorithm Framework
For clarity, we summarize the procedure of RIFT within our AV-centric closed-loop simulation framework in Algorithm 1. The planning model initializes from the IL pre-trained checkpoint provided by Pluto official codebase [^1] https://github.com/jchengai/pluto, followed by closed-loop RL fine-tuning within the CARLA simulator carla to generate realistic and controllable traffic scenarios.
Algorithm 1 Procedure for RIFT in the AV-Centric Closed-Loop Simulation Framework.
1:Input: IL pre-trained planning model $\pi_{\theta_{\text{init}}}$, buffer $\mathcal{D}$ $\triangleright$ IL pre-training (nuPlan caesar2021nuplan )
2:planning model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$
3:for $\text{iteraction}=1,\dots,I$ do $\triangleright$ Closed-loop RL fine-tuning (CARLA carla )
4:  Update the old planning model $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
@@
Compute simulated rollouts $\{\widetilde{\tau}_{i}\}_{i=1}^{G}$ from $\{\tau_{i}\}_{i=1}^{G}$ $\triangleright$ Forward simulation
@@
Compute reward $\{r_{i}\}_{i=1}^{G}$, advantage $\{\hat{A}_{i}\}_{i=1}^{G}$ for each $\widetilde{\tau}_{i}$ with Equation˜6
10:
@@
11:   end for
12:  end while
13:  for RIFT $\text{iteration}=1,\dots,\mu$ do $\triangleright$ Policy fine-tuning
14:
@@
15:
Update model $\pi_{\theta}$ by maximizing the RIFT objective (Equation˜8)
@@
17:end for
18:Output: RL fine-tuned planning model